# Analysis of static and dynamic batching algorithms for graph neural networks

### v2

@@ Appendix A Batch and pad algorithms
For completeness, we provide pseudocode for the batch algorithm mentioned in Section 3. The pseudocode provided is a simplified version of the method found in Jraph, although the TF-GNN, ptgnn, and pyTorch libraries all perform something similar. The basic principle to create a single disconnected graph that contains all of the information of the individual graphs in the batch. The batch method in Algorithm 4, works by creating six empty lists to store information about the batch. It creates a list to collect information about the individual graphs. The lists collect node/edge feature vectors, number of node/edges, and sender/receivers node indices. It then loops over the input graphs list, and concatenates the graph’s node features, edge features, number of nodes/edges and sender/receiver indices into the respective batch lists. Finally, the method creates a new graph object where the constructor is fed the concatenated lists. Thereby, from a list of graphs, we have created one super-graph containing all of the information about the smaller graphs. The individual graphs can be rebuilt from the super-graph if desired. Note that the sender and receiver node indices need to be offset by a running counter of the number of nodes already added into the batch.
@@
[…]
@@
The static algorithm’s padding method, pad-nearest-power-2, mentioned in Section 3 works by adding a fake dummy graph to the list of graphs. It adds a graph object with the number of nodes and edges being equal to the number of nodes/edges required to bring the sum of all the nodes/edges in the list of graphs to the nearest power of two. The algorithm is shown in pseudocode in Algorithm 5. When running the static-64 algorithm the padding is done to the nearest multipe of 64.
@@
Appendix B Timing experiments setup
@@
Appendix C Mean versus median in timing measurements
For each timing experiment ten experiments are run and either the mean or median is taken. The median timing results for the MPEU model after running one hundred thousand steps is seen in Fig. 8. The mean results for the same experiments are shown in 9. The difference in the two figures shows the importance of the number of recompilations in the gradient-update method. This is because the mean is affected by outlier measurements, such as recompilations, while the median is not. The fact that for one hundred thousand training steps, the static-64 algorithm has the best median performance but not the best mean performance is due to the number of recompilations which is highest for this algorithm.
Figure 8: Median batching (top) and median gradient-update times (middle), and median combined time (bottom) for varying batch sizes on the AFLOW (left) and QM9 (right) data using the MPEU model. For each datapoint, ten iterations of one hundred training steps are run.
@@ Appendix C Mean versus median in timing measurements
The mean and median timing results for two million steps are shown for the MPEU model in Fig. 10 and in Fig. 11 respectively. We see the mean static-64 gradient-update step times are shifted higher than the median results, showing that the recompilations still affect the mean results. However, the rankings of the mean combined times are the same for the median combined times across algorithms, suggesting for longer training time, in this case two million steps, the effect of recompilations is no longer as significant as we saw earlier for one hundred thousand steps.
@@
Appendix D CPU only timing results
@@
Appendix E Test metric results
@@
Appendix F Static-constant batching
The TF-GNN library tutorial suggests running the static batching algorithm with a constant padding target. As described in Section 3, the algorithm first iterates through all the graphs in the dataset and saves the maximum number of graphs/edges seen in a graph and uses this as a padding target. We evaluated this method on a subset of batch sizes, dataset and model combinations, and compared it to the static-$2^{N}$, static-64 and dynamic algorithms. The results are shown in Table 1. Our results show that the static-constant algorithm performs poorly in comparison to the other algorithms, and as a result it is not used in the main text.
@@ Appendix F Static-constant batching
Appendix G Effect on rankings of the number of training steps
The effect of the number of training steps run on the rankings of the fastest algorithms is depicted in Fig. 15. We can clearly see that for a smaller number of training steps, the dynamic batching algorithm is faster than the static-64 algorithm which is due to the number of recompilations that happen during the gradient-update step. For more training steps, this effect is subdued and the static-64 algorithm is faster.
Figure 15: Running average combined time (gradient-update step and batching) as a function of the total number of training steps run. Here only a single iteration is run for the batch size 32, MPEU model and QM9 dataset for both the dynamic, static-64 and static-$2^{N}$ algorithms.
Appendix H Student t-tests on test metrics
The student t-test was performed in a pairwise fashion between two batching algorithms upon their distribution of final test RMSE values. For the SchNet model on the QM9 dataset for batch size 64, a heatmap of the t-test values is shown in Fig. 16 and the associated p-values in Fig. 17.
Figure 16: Heatmap of pairwise student t-test values on the distribution of test RMSE values for the SchNet model, QM9 data and a batch size of 64.
Figure 17: Heatmap of p-values from the pairwise student t-test on the distribution of test RMSE values for the SchNet model, QM9 data and a batch size of 64.

### v3

@@ Appendix B Batch and pad algorithms
For completeness, we provide pseudocode for the batch algorithm mentioned in Section 3. The pseudocode provided is a simplified version of the method found in Jraph, although the TF-GNN, ptgnn, and pyTorch libraries all perform something similar. The basic principle to create a single disconnected graph that contains all of the information of the individual graphs in the batch. The batch method in Algorithm 2, works by creating six empty lists to store information about the batch. It creates a list to collect information about the individual graphs. The lists collect node/edge feature vectors, number of node/edges, and sender/receiver node indices. It then loops over the input graphs list, and concatenates the graph’s node features, edge features, number of nodes/edges and sender/receiver indices into the respective batch lists. Finally, the method creates a new graph object where the constructor is fed the concatenated lists. Thereby, from a list of graphs, we have created one super-graph containing all of the information about the smaller graphs. The individual graphs can be rebuilt from the super-graph if desired. Note that the sender and receiver node indices need to be offset by a running counter of the number of nodes already added into the batch.
@@
  for $i=0$ to $\text{len(g.senders})$ do
   $l_{r}\text{.append(g.receivers[i] + o)}$
   $l_{s}\text{.append(g.senders[i] + o) }$
@@
The static algorithm’s padding method, pad-nearest-power-2, mentioned in Section 3 works by adding a fake dummy graph to the list of graphs. It adds a graph object with the number of nodes and edges being equal to the number of nodes/edges required to bring the sum of all the nodes/edges in the list of graphs to the nearest power of two. The algorithm is shown in pseudocode in Algorithm 3. When running the static-64 algorithm the padding is done to the nearest multiple of 64.
@@
The static batching algorithm is shown in Algorithm 4. The padding budget estimation is shown in the Algorithm 5.
Algorithm 4 Static batching
Input: batch size $B_{G}$, training graphs $G$, start point $s$
$A\leftarrow[]$ {Accumulated list of graphs}
$n_{n}\leftarrow 0$ {Accumulated # of nodes}
$n_{e}\leftarrow 0$ {Accumulated # of edges}
for $i=s$ to $s+B_{G}$ do
  $A\leftarrow\text{$A$.append($G[i]$)}$
  $n$ += $_{n}\text{ += }\text{$G[i]$.num-nodes}$
  $n_{e}\text{ += }\text{$G[i]$.num-edges}$
  if $len(A)==(B_{G}-1)$ then
   return batch(pad-nearest-power-2(AA, $n_{n}$, $n_{e}$))
  end if
end for
Algorithm 5 Estimate dynamic batching padding target
Input: subset size NN, list of graphs $G$, batch size $B_{G}$
$n\leftarrow 0$ {number of nodes in N graphs}
$e\leftarrow 0$ {number of edges in N graphs}
$g\leftarrow G[1:N]$ {Grab N graphs from training data}
for $i=1$ to $i=N$ do
  $n$ += G[i].G[i].num-nodes
  $e$ += G[i].G[i].num-edges
end for
$B_{n}\leftarrow\text{next-multiple-64}(n*B_{G}/N)$ {padding target}
$B_{e}\leftarrow\text{next-multiple-64}(e*B_{G}/N)$
return $B_{n}$, $B_{e}$, $B_{G}$
Appendix C Timing experiments setup
@@
Appendix D Mean versus median in timing measurements
For each timing experiment ten experiments are run and either the mean or median is taken. The median timing results for the MPEU model after running one hundred thousand steps is seen in Fig. 6. The mean results for the same experiments are shown in Fig. 7. The difference in the two figures shows the importance of the number of recompilations in the gradient-update method. This is because the mean is affected by outlier measurements, such as recompilations, while the median is not. The fact that for one hundred thousand training steps, the static-64 algorithm has the best median performance but not the best mean performance is due to the number of recompilations which is highest for this algorithm.
Figure 6: Median batching (top) and median gradient-update times (middle), and median combined time (bottom) for varying batch sizes on the AFLOW (left) and QM9 (right) data using the MPEU model. For each datapoint, ten iterations of one hundred thousand training steps are run.
@@ Appendix D Mean versus median in timing measurements
The mean and median timing results for two million steps are shown for the MPEU model in Fig. 8 and in Fig. 9 respectively. We see the mean static-64 gradient-update step times are shifted higher than the median results, showing that the recompilations still affect the mean results. However, the rankings of the mean combined times are the same for the median combined times across algorithms, suggesting that for longer training time the effect of recompilations is no longer as significant as we saw earlier for one hundred thousand steps.
@@
Appendix E CPU only timing results
@@
Appendix F Multi-GPU behavior
We used JAX parallel map (pmap) function to run the batching algorithms on a multi-GPUs of four GPUs. In this case, we saw that the static-64 and static-$2^{n}$ algorithms did not work out of the box with pmap since the algorithm expects the batches sent to each GPU to be the same size. For the AFLOW dataset, we saw roughly that the four batches had different shapes after padding when using the static-$2^{n}$ algorithm roughly 90% of the time, and close to 99% when using the static-64 algorithm. It’s possible to alter the static-64 and static-$2^{N}$} algorithm to pad each batch of data (e.g. for four GPUs there are four batches) using the largest number of nodes and edges across all batches. The results of such tests, however, will be highly implementation dependent, more-so than in the single-GPU case. To simplify the comparison of the algorithms, we evaluate the algorithms only on single-GPU setups and note that the dynamic algorithm is easily ported to a multi-GPU setup using JAX’s pmap function.
Appendix G Test metric results
@@
Appendix H Static-constant batching
The TF-GNN library tutorial suggests running the static batching algorithm with a constant padding target. As described in Section 3, the algorithm first iterates through all the graphs in the dataset and saves the maximum number of nodes/edges seen in a graph and uses this as a padding target. We evaluated this method on a subset of batch sizes, dataset and model combinations, and compared it to the $2^{N}, static$-64 and dynamic algorithms. The results are shown in Table 1. Our results show that the static-constant algorithm performs poorly in comparison to the other algorithms, and as a result it is not used in the main text.
@@ Appendix H Static-constant batching
static-constant AFLOW PaiNN 16 0.83 46.56 47.39
static-$2^{N}$ AFLOW PaiNN 16 0.45 3.82 4.27
static-6464 AFLOW PaiNN 16 0.46 3.88 4.34
dynamic AFLOW PaiNN 16 0.59 3.20 3.79
@@
static-constant QM9 PaiNN 16 0.94 11.67 12.61
static-$2^{N}$ QM9 PaiNN 16 0.71 6.43 7.14
static-6464 QM9 PaiNN 16 0.50 3.30 3.80
dynamic QM9 PaiNN 16 0.68 3.33 4.01
Appendix I Student’s t-tests on test metrics
The Student’s t-test was performed in a pairwise fashion between two batching algorithms upon their distribution of final test RMSE values. For the SchNet model on the QM9 dataset for batch size 64, a heatmap of the t-test values is shown in Fig. 13 and the associated p-values in Fig. 14.
Figure 13: Heatmap of pairwise Student’s t-test values on the distribution of test RMSE values for the SchNet model, QM9 data and a batch size of 64.
Figure 14: Heatmap of p-values from the pairwise Student’s t-test on the distribution of test RMSE values for the SchNet model, QM9 data and a batch size of 64.